Primo car rental agency charges $25 per day plus $0.30 per mile. Ultimo car rental agency charges $15 per day plus $1.00 per mile. Find the daily mileage for which the Ultimo charge is three times the Primo charge.
Answer by Boreal(15235)

3(25+0.3x)=15+1x, the two functions are linear.
75+0.9x=15+1x
60=0.1x
x=600 miles
then Primo's charge is $25+180=$205 and Ultimo's is $615